When 2.00 mol of SO2Cl2 is placed in a 2.00-L flask at 302 K , 58 % of the SO2Cl2 decomposes to SO2 and Cl2:
SO2Cl2(g)⇌SO2(g)+Cl2(g)

B) Calculate Kc for 2.00 mol of SO2Cl2 in a 13.00 −L vessel at 302 K .

Express your answer to two significant figures.
Answer #1

SO2Cl2 molarity = 2/2 =1

SO2Cl2 (g) <----------------------> SO2 (g) + Cl2 (g)

1                                                   0               0            ----------------------initial

1-x                                               x               x -------------------------------equilibrium

1-0.58 0.568 0.58

0.42 0.58 0.58

decomposition

x = 58 %

   = 58/100

    = 0.58

Kc = [SO2][SO2]/ [SO2Cl2]

Kc = 0.58 x 0.58 / 0.42

Kc = 0.801